# Nonadiabatic corrections to electric quadrupole transition rates in H2

### v1

@@ 1 Introduction
In this work, we derive a formula for the leading nonadiabatic correction to rates of electric quadrupole transitions in H2. This nonadiabatic correction can be represented in terms of a single function $D^{(1)}(R),$ which can be added to the BO quadrupole moment function $D^{(0)}(R)$, calculated by Wolniewicz in Ref. [9], recalculated by Komasa [10] using ECG functions, and significantly improved here using an enhanced version of H2SOLV code [11], which employs the Kołos-Wolniewicz (KW) functions. Apart from many applications in astrophysics, these accurate rates can be used in the primary thermometry for temperatures as low as 10 K [4, 12, 13].
@@ 2 Nonadiabatic perturbation theory
where the summation index $a$ enumerates all electrons, and $A$ and $B$ refer to nuclei. To derive formulae for nonadiabatic effects, one must fix the reference frame. We start with the laboratory frame $\{\vec{r}_{A},\,\vec{r}_{B},\,\vec{r}_{a}\}$, and subsequently change variables to $\{\vec{R},\,\vec{R}_{G},\,\vec{x}_{a}\}$ according to
@@
Next, the angular momentum operator $\vec{J}$, for states with vanishing total momentum, is
$\vec{J}=\sum_{a}(\vec{r}_{a}-\vec{R}_{G})\times\vec{p}_{a}+(\vec{r}_{A}-\vec{R}_{G})\times\vec{p}_{A}+(\vec{r}_{B}-\vec{R}_{G})\times\vec{p}_{B},$ (12)
and in new variables
$\vec{J}=\sum_{a}\vec{x}_{a}\times\vec{q}_{a}+\vec{R}\times\vec{P}\equiv\vec{J}_{\rm el}+\vec{J}_{\rm n}$ (13)
it can also be split into electronic $\vec{J}_{\rm el}$ and nuclear $\vec{J}_{\rm n}$ parts.
@@ 3 Electric quadrupole moment
where the summation index $\alpha$ refers to both electrons and nuclei, and $\vec{r}_{\alpha M}$ are particle positions with respect to the center of mass, $\vec{r}_{M}$,
@@
Its matrix element on the $\phi_{\mathrm{el}}$ state in the BO-approximation is
@@
$Q_{3}(R)=$ $\ -\frac{2\,i}{R^{3}}\langle\vec{\psi}_{\Pi}|\vec{R}\times\vec{J}_{\mathrm{el}}|\phi_{\mathrm{el}}\rangle,$ (61)
@@ 4 $R$-derivatives
Let us further consider the first-order $R$-derivative of $\phi$
$\partial_{R}\phi=\sum_{k}\left(\psi_{k}\,\partial_{R}v_{k}+v_{k}\,\partial_{R}\psi_{k}\right)\,.$ (70)
@@
and add it to the nonrelativistic Hamiltonian $H$. So that $Q_{2}(R)$ can be obtained by a differentiation in $\xi$,
$Q_{2}=$ $\ \frac{1}{2}\,\frac{d}{d\,\xi}\biggr|_{\xi=0}\,\langle\nabla^{k}_{R}\phi_{\mathrm{el}}|\nabla^{k}_{R}\phi_{\mathrm{el}}\rangle_{\xi}\,.$ (77)
@@ 5 Numerical calculations
We employ a variational approach to obtain $\phi_{\mathrm{el}}$ for the ground electronic $\Sigma^{+}$ state and utilize explicitly correlated exponential functions with polynomial dependence on interparticle distances of the form [15]
@@
where $r_{\perp}^{i}=r^{i}-n^{i}\,\vec{n}\,\vec{r}$, $\hat{S}^{\pm}_{AB}=1\pm P_{AB}$ and $P_{AB}$ permutes the nuclei $A$ and $B$, $\hat{S}^{\pm}_{12}=1\pm P_{12}$ and $P_{12}$ interchange the two electrons, and appropriate $\pm$ signs are chosen to fulfill the symmetry criteria for gerade/ungerade and singlet/triplet states. By solving the secular equation one obtains linear coefficients $v_{\{n\}}$. Such a form of wavefunction expansion is commonly referred to as the Kołos-Wolniewicz basis [15]. In our calculations we use two specific cases of the KW basis, namely the symmetric James-Coolidge (JC) basis ($y=x=0,\alpha=u=w$) for $R\leq 10$ and symmetric Heitler-London (HL) basis ($\alpha=-y=x=u=w$) for $R\geq 10$. Consequently, the resulting parametrization of our trial basis is extremely compact – just a single nonlinear parameter $\alpha$ and the integer shell constraint $\Omega$.
@@
The efficient recursive evaluation of $f$-integrals is described in Refs. [11, 16].
@@
which are obtained by solving a corresponding linear equation in the basis of $\Psi_{\Sigma^{+}}$ and $\Psi_{\Pi}$ functions, respectively. Those bases for intermediate states carry exactly the same set of nonlinear parameters as the external basis – optimized for the energy of the ground electronic state.
@@ 6 Transition rates
The radial matrix element in Eq. (91) is calculated with $D(R)$ given by Eq. (55) and with nuclear wavefunctions $\chi$ being a solution of the radial equation with the inclusion of the diagonal adiabatic correction. We note that the $A_{\mathrm{E2}}$ rate is sensitive to this transition energy; therefore, we take values from the H2Spectre code [19], which are accurate to about 1 MHz.
Our numerical results are in agreement with those of Wolniewicz [9] and Komasa [10], but should be more accurate due to the inclusion of the leading nonadiabatic correction $D^{(1)}$. In fact, the E2 transition rates’ uncertainty is now dominated by the unknown relativistic corrections, which we estimated by $\alpha^{2}$ times $A_{\mathrm{E2}}$. For the M1 transition rates, we use Eq. (13) from Ref. [10], and our results are also in agreement with that work.
In Table 6 we present the transition rates for the Q ($J^{\prime\prime}=J^{\prime})$, S ($J^{\prime\prime}=J^{\prime}-2$), and O ($J^{\prime\prime}=J^{\prime}+2$) branches of the fundamental ($\nu=1\rightarrow 0$) band of H2.
@@
Transition rates for the Q ($J^{\prime\prime}=J^{\prime})$, S ($J^{\prime\prime}=J^{\prime}-2$), and O ($J^{\prime\prime}=J^{\prime}+2$) branches of the fundamental ($\nu=1\rightarrow 0$) band of H2 in units of $10^{-8}$ s-1. $A_{\mathrm{M1}}$ is the magnetic dipole transition rate with the uncertainty due to unknown finite-nuclear-mass corrections estimated as $\sim 2\frac{m_{e}}{m_{\mathrm{n}}}A_{\mathrm{M1}}$. $\delta A_{X}\equiv A_{X}^{(1)}-A_{X}^{(0)}$, where $A_{X}^{(0)}$ is the E2 transition rate in the BO approximation ($D^{(0)}$ operator and BO nuclear wavefunction) with $X=Q,S,O$, for Q-, S-, and O-branch, respectively; whereas $A_{X}\equiv A_{X}^{(1)}$ incorporates the mass-corrected quadrupole operator ($D^{(1)}$) with the adiabatically corrected nuclear wavefunction. The E2 rates’ uncertainty is now dominated by the unknown relativistic corrections, estimated by $\sim\alpha^{2}A_{\mathrm{E2}}\approx 0.5\cdot 10^{-4}A_{\mathrm{E2}}$. Physical constants are from Ref. [17]. $J^{\prime}$ $A_{\mathrm{M1}}$ $A_{Q}$ $10^{2}\cdot\,\delta A_{Q}/A_{Q}$ $A_{S}$ $10^{2}\cdot\,\delta A_{S}/A_{S}$ $A_{O}$ $10^{2}\cdot\,\delta A_{O}/A_{O}$ 0 $0.0$ $0.0$ - $0.0$ - 85.601 285.601\,2 $0.43$ 1 0.071 2630.071\,263 43.018 5843.018\,58 $0.44$ $0.0$ - 42.355 5742.355\,57 $0.43$ 2 0.213 8560.213\,856 30.394 7530.394\,75 $0.44$ 25.345 5725.345\,57 $0.45$ 29.048 5029.048\,50 $0.42$ 3 0.427 9000.427\,900 27.907 3227.907\,32 $0.44$ 34.821 5834.821\,58 $0.47$ 20.885 1020.885\,10 $0.42$ 4 0.713 5350.713\,535 26.592 3526.592\,35 $0.44$ 39.957 9839.957\,98 $0.48$ 15.022 8715.022\,87 $0.42$ 5 1.070 8881.070\,888 25.537 0825.537\,08 $0.45$ 42.228 1642.228\,16 $0.50$ 10.645 7710.645\,77 $0.41$ 6 1.500 0021.500\,002 24.520 3724.520\,37 $0.45$ 42.056 3842.056\,38 $0.52$ 7.379 7407.379\,740 $0.41$ 7 2.000 8582.000\,858 23.471 4823.471\,48 $0.45$ 39.717 0939.717\,09 $0.55$ 4.983 0364.983\,036 $0.41$ 8 2.573 1762.573\,176 22.367 3322.367\,33 $0.46$ 35.542 9035.542\,90 $0.59$ 3.266 3923.266\,392 $0.41$ 9 3.216 4973.216\,497 21.203 8821.203\,88 $0.46$ 29.974 9529.974\,95 $0.64$ 2.071 5602.071\,560 $0.41$ 10 3.930 0603.930\,060 19.985 8819.985\,88 $0.47$ 23.555 9423.555\,94 $0.71$ 1.266 1241.266\,124 $0.41$ 11 4.712 7474.712\,747 18.722 6418.722\,64 $0.47$ 16.897 1516.897\,15 $0.82$ 0.742 08040.742\,0804 $0.41$ 12 5.563 0335.563\,033 17.425 8817.425\,88 $0.48$ 10.634 7710.634\,77 $1.00$ 0.414 35830.414\,3583 $0.41$ 13 6.478 9076.478\,907 16.108 5216.108\,52 $0.49$ 5.384 7765.384\,776 $1.34$ 0.218 46280.218\,4628 $0.42$ 14 7.457 7977.457\,797 14.783 9414.783\,94 $0.50$ 1.702 8941.702\,894 $2.26$ 0.107 40670.107\,4067 $0.42$ 15 8.496 4748.496\,474 13.465 4213.465\,42 $0.51$ 0.053 15050.053\,1505 $11.79$ 0.048 36040.048\,3604 $0.42$ 16 9.590 9359.590\,935 12.165 8012.165\,80 $0.52$ 0.786 64110.786\,6411 2.98-2.98 0.019 40460.019\,4046 $0.42$ 17 10.736 2610.736\,26 10.897 1510.897\,15 $0.53$ 4.130 4134.130\,413 1.21-1.21 0.006 64200.006\,6420 $0.43$ 18 11.926 4311.926\,43 9.670 589.670\,58 $0.55$ 10.185 2610.185\,26 0.71-0.71 0.001 79740.001\,7974 $0.43$ 19 13.154 1213.154\,12 8.496 088.496\,08 $0.56$ 18.930 5718.930\,57 0.48-0.48 0.000 33090.000\,3309 $0.43$ 20 14.410 3714.410\,37 7.382 427.382\,42 $0.58$ 30.234 2330.234\,23 0.35-0.35 0.000 02860.000\,0286 $0.44$ 21 15.684 3015.684\,30 6.337 126.337\,12 $0.60$ 43.865 4243.865\,42 0.27-0.27 0.000 000 2700.000\,000\,270 $0.44$ 22 16.962 6016.962\,60 5.366 345.366\,34 $0.63$ 59.508 7159.508\,71 0.21-0.21 23 18.228 9818.228\,98 4.474 974.474\,97 $0.66$ 76.777 7376.777\,73 0.17-0.17 24 19.463 5119.463\,51 3.666 563.666\,56 $0.69$ 95.227 495.227\,4 0.14-0.14 25 20.641 6220.641\,62 2.943 402.943\,40 $0.73$ 114.363 4114.363\,4 0.12-0.12 26 21.732 9621.732\,96 2.306 502.306\,50 $0.78$ 133.648 6133.648\,6 0.11-0.11 27 22.699 7522.699\,75 1.755 611.755\,61 $0.84$ 152.503 8152.503\,8 0.10-0.10 28 23.494 6723.494\,67 1.289 291.289\,29 $0.92$ 170.303 9170.303\,9 0.09-0.09 29 24.057 7924.057\,79 0.904 820.904\,82 $1.01$ 186.363 0186.363\,0 0.09-0.09 30 24.311 9824.311\,98 0.598 280.598\,28 $1.14$ 199.905 3199.905\,3 0.08-0.08
@@
In the Q-branch, we note that except for $J^{\prime}=1$, the magnetic dipole $A_{\mathrm{M1}}$ rate exceeds the nonadiabatic corrections to $A_{Q}$, significantly grows with $J^{\prime}$, and starting from $J^{\prime}=18$ exceeds the total $A_{Q}$. In the O-branch, the relative nonadiabatic correction is nearly constant and amounts to $\approx 0.4\%$. Whereas for the S-branch, the relative nonadiabatic correction varies strongly, starting from $\approx 0.45\%$ at $J^{\prime}=2$, achieving a maximum of 12% at $J^{\prime}=15$ due to the smallness of the radial matrix element. This behaviour might be interesting to verify experimentally.

### v2

@@ 1 Introduction
In this work, we derive a formula for the leading nonadiabatic correction to rates of electric quadrupole transitions in H2. This nonadiabatic correction can be represented in terms of a single function $D^{(1)}(R),$ which can be added to the BO quadrupole moment function $D^{(0)}(R)$, calculated by Wolniewicz in Ref. [10], recalculated by Komasa [11] using ECG functions, and significantly improved here using an enhanced version of H2SOLV code [12], which employs the Kołos-Wolniewicz (KW) functions. Apart from many applications in astrophysics, these accurate rates can be used in the primary thermometry for temperatures as low as 10 K [4, 13, 14, 5]. In particular Authors of Ref. [14] propose to measure the ratio of two transition rates from the same vibrational band to achieve higher accuracy for temperature due to cancellation of experimental and theoretical uncertainties. Hydrogen molecule is the best for this purpose, because transition rates can be calculated very accurately from first principles, and this work is the first step toward this accurate termometry.
@@ 2 Nonadiabatic perturbation theory
where the summation index $a$ enumerates all electrons, and $A$ and $B$ refer to nuclei, and atomic units are used throughout the paper. To derive formulae for nonadiabatic effects, one must fix the reference frame. We start with the laboratory frame $\{\vec{r}_{A},\,\vec{r}_{B},\,\vec{r}_{a}\}$, and subsequently change variables to $\{\vec{R},\,\vec{R}_{G},\,\vec{x}_{a}\}$ according to
@@
Next, the total angular momentum operator $\vec{J}$ does not depend on the reference point, because we assumed the vanishing total momentum for molecular states. Therefore we chose for convenience $\vec{R}_{G}$ as a reference point and $J$ becomes
$\vec{J}=\sum_{a}(\vec{r}_{a}-\vec{R}_{G})\times\vec{p}_{a}+(\vec{r}_{A}-\vec{R}_{G})\times\vec{p}_{A}+(\vec{r}_{B}-\vec{R}_{G})\times\vec{p}_{B}\,.$ (12)
In new variables
$\vec{J}=\sum_{a}\vec{x}_{a}\times\vec{q}_{a}+\vec{R}\times\vec{P}\equiv\vec{J}_{\rm el}+\vec{J}_{\rm n},$ (13)
$J$ can also be split into electronic $\vec{J}_{\rm el}$ and nuclear $\vec{J}_{\rm n}$ parts.
@@ 3 Electric quadrupole moment
where the summation index $\alpha$ refers to both electrons and nuclei, $e_{\alpha}$ refers to the electric charge of the $\alpha$-th particle, and $\vec{r}_{\alpha M}$ are particle positions with respect to the center of mass, $\vec{r}_{M}$,
@@
Its matrix element on the $\phi_{\mathrm{el}}$ state in the BO approximation is
@@
$Q_{3}(R)=$ $\ -\frac{2\,i}{R^{3}}\langle\vec{\phi}_{\Pi}|\vec{R}\times\vec{J}_{\mathrm{el}}|\phi_{\mathrm{el}}\rangle,$ (61)
@@ 4 $R$-derivatives
Let us further consider the first-order $R$-derivative of $\phi_{\mathrm{el}}$
$\partial_{R}\phi_{\mathrm{el}}=\sum_{k}\left(\psi_{k}\,\partial_{R}v_{k}+v_{k}\,\partial_{R}\psi_{k}\right)\,.$ (70)
@@
and add it to the nonrelativistic Hamiltonian $H_{\mathrm{el}}$. So that $Q_{2}(R)$ can be obtained by a differentiation in $\xi$,
$Q_{2}=$ $\ \frac{1}{2}\,\frac{d}{d\,\xi}\biggr|_{\xi=0}\,\langle\nabla^{k}_{R}\phi_{\mathrm{el}}|\nabla^{k}_{R}\phi_{\mathrm{el}}\rangle_{\xi}\,$, (77)
where the subscript $\xi$ in $\langle\ldots\rangle_{\xi}$ stands for the $\xi$-dependence of $\phi_{\mathrm{el}}$ due to $\xi$-dependence of the Hamiltonian in Eq. (76). The methods for calculation of $Q_{2}$ are described in the next Section and in Appendix A.
@@ 5 Numerical calculations
Following Ref. [16] we employ a variational approach to obtain $\phi_{\mathrm{el}}$ for the ground electronic $\Sigma^{+}$ state and utilize explicitly correlated exponential functions with polynomial dependence on interparticle distances of the form [17]
@@
where $r_{\perp}^{i}=r^{i}-n^{i}\,\vec{n}\,\vec{r}$, $\hat{S}^{\pm}_{AB}=1\pm P_{AB}$ and $P_{AB}$ permutes the nuclei $A$ and $B$, $\hat{S}^{\pm}_{12}=1\pm P_{12}$ and $P_{12}$ interchange the two electrons, and appropriate $\pm$ signs are chosen to fulfill the symmetry criteria for gerade/ungerade and singlet/triplet states. By solving the secular equation one obtains linear coefficients $v_{\{n\}}$. Such a form of wavefunction expansion is commonly referred to as the Kołos-Wolniewicz basis [17]. In our calculations we use two specific cases of the KW basis, namely the symmetric James-Coolidge (JC) basis ($y=x=0,\alpha=u=w$) for $R\leq 10$ au and symmetric Heitler-London (HL) basis ($\alpha=-y=x=u=w$) for $R\geq 10$ au. Consequently, the resulting parametrization of our trial basis is extremely compact – just a single nonlinear parameter $\alpha$ and the integer shell constraint $\Omega$.
@@
The efficient recursive evaluation of $f$-integrals in arbitrary (extended) precision is described in Refs. [12, 18]. The numerical algebra part of the calculations is performed using parallel HSL_mp54 Cholesky solver [19, 20] adapted to quad-double arithmetic precision (64 decimal digits) [21].
@@
which are obtained by solving a corresponding linear equation in the basis of $\Psi_{\Sigma^{+}}$ and $\Psi_{\Pi}$ functions, respectively. Those bases for intermediate states carry exactly the same set of nonlinear parameters as the external basis – optimized for the energy of the ground electronic state. This considerably reduces the number of integrals that have to be computed, and for the case of $|\phi_{\Sigma^{\!+}}\!\rangle$ allows us to reuse the basis already constructed for the external (ground) state, hence we have $\Omega_{\Sigma^{+}}=\Omega$. Whereas for $|\phi_{\Pi}^{k}\rangle$ we choose $\Omega_{\Pi}=\Omega-2$, which curiously, due to different symmetry restrictions for $\Sigma_{g}^{+}$ and $\Pi_{g}$, results in almost the same basis sizes, both in JC and HL cases.
@@ 6 Transition rates
The radial matrix element in Eq. (91) is calculated with $D(R)$ given by Eq. (55) and with nuclear wavefunctions $\chi$ being a solution of the radial equation with the inclusion of the diagonal adiabatic correction. We note that the $A_{\mathrm{E2}}$ rate is sensitive to this transition energy; therefore, we take values from the H2Spectre code [24], which are accurate to about 1 MHz. Our numerical results are in agreement with those of Wolniewicz [10] and Komasa [11], but should be more accurate due to the inclusion of the leading nonadiabatic correction $D^{(1)}$. In fact, the E2 transition rates’ uncertainty is now dominated by the unknown relativistic corrections, which we estimated by $\alpha^{2}$ times $A_{\mathrm{E2}}$. We note, that magnetic dipole (M1) transitions occur at the same wavelengths as the electric quadrupole (E2) transitions. Consequently, in the spontaneous emission the total transition probability is the sum of the separate M1 and E2 rates. Therefore, whenever transition rates in the Q-branch ($J^{\prime\prime}=J^{\prime}$) are of interest, one has to include also the M1 channel. For the M1 transition rates in Table 6, we use Eq. (13) from Ref. [11], and our results are in a good agreement.
In Table 6, apart from $A_{\mathrm{M1}}$, we present the $A_{\mathrm{E2}}^{Q}$, $A_{\mathrm{E2}}^{S}$, and $A_{\mathrm{E2}}^{O}$ transition rates for the Q ($J^{\prime\prime}=J^{\prime})$, S ($J^{\prime\prime}=J^{\prime}-2$), and O ($J^{\prime\prime}=J^{\prime}+2$) branches of the fundamental ($\nu=1\rightarrow 0$) band of H2, respectively.
@@
Transition rates for the Q ($J^{\prime\prime}=J^{\prime})$, S ($J^{\prime\prime}=J^{\prime}-2$), and O ($J^{\prime\prime}=J^{\prime}+2$) branches of the fundamental ($\nu=1\rightarrow 0$) band of H2 in units of $10^{-8}$ s-1. $A_{\mathrm{M1}}$ is the magnetic dipole transition rate with the uncertainty due to unknown finite-nuclear-mass corrections estimated as $\sim 2\frac{m_{e}}{m_{\mathrm{n}}}A_{\mathrm{M1}}$. $\delta A_{\mathrm{E2}}^{X}$ is the difference in the E2 transition rate between those calculated with the mass-corrected quadrupole operator ($D^{(1)}$) with the adiabatically corrected nuclear wavefunction with $X=Q,S,O$, for Q-, S-, and O-branch, respectively; and the E2 rate in the BO approximation ($D^{(0)}$ operator and BO nuclear wavefunction). The E2 rates’ uncertainty is now dominated by the unknown relativistic corrections, estimated by $\sim\alpha^{2}A_{\mathrm{E2}}\approx 0.5\cdot 10^{-4}A_{\mathrm{E2}}$. Physical constants are from Ref. [22]. $J^{\prime}$ $A_{\mathrm{M1}}$ $A_{\mathrm{E2}}^{Q}$ $10^{2}\cdot\,\delta A_{\mathrm{E2}}^{Q}/A_{\mathrm{E2}}^{Q}$ $A_{\mathrm{E2}}^{S}$ $10^{2}\cdot\,\delta A_{\mathrm{E2}}^{S}/A_{\mathrm{E2}}^{S}$ $A_{\mathrm{E2}}^{O}$ $10^{2}\cdot\,\delta A_{\mathrm{E2}}^{O}/A_{\mathrm{E2}}^{O}$ 0 $0.0$ $0.0$ - $0.0$ - 85.601 285.601\,2 $0.43$ 1 0.071 2630.071\,263 43.018 5843.018\,58 $0.44$ $0.0$ - 42.355 5742.355\,57 $0.43$ 2 0.213 8560.213\,856 30.394 7530.394\,75 $0.44$ 25.345 5725.345\,57 $0.45$ 29.048 5029.048\,50 $0.42$ 3 0.427 9000.427\,900 27.907 3227.907\,32 $0.44$ 34.821 5834.821\,58 $0.47$ 20.885 1020.885\,10 $0.42$ 4 0.713 5350.713\,535 26.592 3526.592\,35 $0.44$ 39.957 9839.957\,98 $0.48$ 15.022 8715.022\,87 $0.42$ 5 1.070 8881.070\,888 25.537 0825.537\,08 $0.45$ 42.228 1642.228\,16 $0.50$ 10.645 7710.645\,77 $0.41$ 6 1.500 0021.500\,002 24.520 3724.520\,37 $0.45$ 42.056 3842.056\,38 $0.52$ 7.379 7407.379\,740 $0.41$ 7 2.000 8582.000\,858 23.471 4823.471\,48 $0.45$ 39.717 0939.717\,09 $0.55$ 4.983 0364.983\,036 $0.41$ 8 2.573 1762.573\,176 22.367 3322.367\,33 $0.46$ 35.542 9035.542\,90 $0.59$ 3.266 3923.266\,392 $0.41$ 9 3.216 4973.216\,497 21.203 8821.203\,88 $0.46$ 29.974 9529.974\,95 $0.64$ 2.071 5602.071\,560 $0.41$ 10 3.930 0603.930\,060 19.985 8819.985\,88 $0.47$ 23.555 9423.555\,94 $0.71$ 1.266 1241.266\,124 $0.41$ 11 4.712 7474.712\,747 18.722 6418.722\,64 $0.47$ 16.897 1516.897\,15 $0.82$ 0.742 08040.742\,0804 $0.41$ 12 5.563 0335.563\,033 17.425 8817.425\,88 $0.48$ 10.634 7710.634\,77 $1.00$ 0.414 35830.414\,3583 $0.41$ 13 6.478 9076.478\,907 16.108 5216.108\,52 $0.49$ 5.384 7765.384\,776 $1.34$ 0.218 46280.218\,4628 $0.42$ 14 7.457 7977.457\,797 14.783 9414.783\,94 $0.50$ 1.702 8941.702\,894 $2.26$ 0.107 40670.107\,4067 $0.42$ 15 8.496 4748.496\,474 13.465 4213.465\,42 $0.51$ 0.053 15050.053\,1505 $11.79$ 0.048 36040.048\,3604 $0.42$ 16 9.590 9359.590\,935 12.165 8012.165\,80 $0.52$ 0.786 64110.786\,6411 2.98-2.98 0.019 40460.019\,4046 $0.42$ 17 10.736 2610.736\,26 10.897 1510.897\,15 $0.53$ 4.130 4134.130\,413 1.21-1.21 0.006 64200.006\,6420 $0.43$ 18 11.926 4311.926\,43 9.670 589.670\,58 $0.55$ 10.185 2610.185\,26 0.71-0.71 0.001 79740.001\,7974 $0.43$ 19 13.154 1213.154\,12 8.496 088.496\,08 $0.56$ 18.930 5718.930\,57 0.48-0.48 0.000 33090.000\,3309 $0.43$ 20 14.410 3714.410\,37 7.382 427.382\,42 $0.58$ 30.234 2330.234\,23 0.35-0.35 0.000 02860.000\,0286 $0.44$ 21 15.684 3015.684\,30 6.337 126.337\,12 $0.60$ 43.865 4243.865\,42 0.27-0.27 0.000 000 2700.000\,000\,270 $0.44$ 22 16.962 6016.962\,60 5.366 345.366\,34 $0.63$ 59.508 7159.508\,71 0.21-0.21 23 18.228 9818.228\,98 4.474 974.474\,97 $0.66$ 76.777 7376.777\,73 0.17-0.17 24 19.463 5119.463\,51 3.666 563.666\,56 $0.69$ 95.227 495.227\,4 0.14-0.14 25 20.641 6220.641\,62 2.943 402.943\,40 $0.73$ 114.363 4114.363\,4 0.12-0.12 26 21.732 9621.732\,96 2.306 502.306\,50 $0.78$ 133.648 6133.648\,6 0.11-0.11 27 22.699 7522.699\,75 1.755 611.755\,61 $0.84$ 152.503 8152.503\,8 0.10-0.10 28 23.494 6723.494\,67 1.289 291.289\,29 $0.92$ 170.303 9170.303\,9 0.09-0.09 29 24.057 7924.057\,79 0.904 820.904\,82 $1.01$ 186.363 0186.363\,0 0.09-0.09 30 24.311 9824.311\,98 0.598 280.598\,28 $1.14$ 199.905 3199.905\,3 0.08-0.08
@@
In the Q-branch, we note that except for $J^{\prime}=1$, the magnetic dipole $A_{\mathrm{M1}}$ rate exceeds the nonadiabatic corrections to $A_{\mathrm{E2}}^{Q}$, significantly grows with $J^{\prime}$, and starting from $J^{\prime}=18$ exceeds the total $A_{\mathrm{E2}}^{Q}$. In the ($v=1\rightarrow 0$) O-branch, the relative nonadiabatic correction is nearly constant and amounts to $\approx 0.4\%$, the rapid drop in the magnitude of the overall E2 rate (as $J^{\prime}$ is increased) in this branch can be attributed to the strong suppression by the $(E_{v^{\prime}J^{\prime}}-E_{v^{\prime\prime}J^{\prime\prime}})^{5}$ factor in Eq. (91) as the energy difference goes to zero around $J^{\prime}=21$. Whereas for the S-branch, the relative nonadiabatic correction varies strongly, starting from $\approx 0.45\%$ at $J^{\prime}=2$, achieving a maximum of 12% at $J^{\prime}=15$ due to the smallness of the radial matrix element. This behaviour might be interesting to verify experimentally.